Yerbury Corp. manufactures construction equipment.

Journalize the entries to record the following selected equity investment transactions completed by Yerbury during a recent year:

Feb. 2 Purchased for cash 1,050 shares of Wong Inc. stock for $53 per share plus a $525 brokerage commission.
Mar. 16 Received dividends of $0.30 per share on Wong Inc. stock.
June 7 Purchased 650 shares of Wong Inc. stock for $60 per share plus a $325 brokerage commission.
July 26 Sold 1,200 shares of Wong Inc. stock for $65 per share less a $600 brokerage commission. Yerbury assumes that the first investments purchased are the first investments sold.
Sept. 25 Received dividends of $0.40 per share on Wong Inc. stock.

In your computations, round per share amounts to two decimal places. When required, round final answers to the nearest dollar. For a compound transaction, if an amount box does not require an entry, leave it blank.

Feb. 2 Investments-Wong Inc. Stock   
Cash
Mar. 16 Cash
Dividend Revenue
June 7 Investments-Wong Inc. Stock
Cash
July 26 Cash
***Can someone explain how to get the answers here --> Gain on Sale of Investments
**--------------------------------------------------> Investments-Wong Inc. Stock
Sept. 25 Cash
Dividend Revenue
Answer #1

D Credit($) A 1 Date Particulars 2-Feb Investments-Wong Inc. stock Cash Debit($) 56175 56175 16-Mar Cash 315 Dividend revenueWorkings:

D А 1 Date 2 43863 Credit($) B Particulars Investments-Wong Inc. stock Cash Debit($) =(1050*53)+525 56175 43906 Cash =1050*0.